Aliyah is preparing to expand her IT consulting company. The current market rate for IT professionals is $58,000 per year. Each employee she hires will also require a computer and equipment that costs $6,000 per employee annually. Hiring more employees means that Aliyah can provide consulting services to more clients each year. Each client Aliyah has will pay her $10,000 per year.

The number of clients Aliyah can take on is dependent on the number of workers she hires and is depicted in the accompanying table.

Use this information to calculate the marginal cost and the marginal benefit of hiring each worker.

Number of workers Clients per year
0 0
1 11
2 20
3 27
4 32

a. The first worker's marginal cost is $__________.

b. The first worker's marginal benefit is $__________.

c. The second worker's marginal cost is $__________.

d. The second worker's marginal benefit is $__________.

e. The third worker's marginal cost is $__________.

f. The third worker's marginal benefit is $__________.

g. The fourth worker's marginal cost is $__________.

h. The fourth worker's marginal benefit is $__________.

i. Using the rational rule to maximize her economic surplus, Aliyah should hire _________ workers.
Answer #1

Answer a : The first worker marginal cost is $64000 per worker anual.

b : The marginal benefit per worker for hiring first worker is $1,10,000.

c : The second worker marginal cost is $64000 per worker : anually.

d: The second worker marginal benefit is $10,000*9= $90,000

e : The third worker marginal cost is $64000 per worker anually.

f: The third worker marginal benefit is $10,000*7= $70,000

g : The fourth worker marginal cost is $64000

h : The fourth worker marginal benefit is $10,000*5=$50,000

I . Aliyah should hired 3 workers as the additional worker marginal benefit is more than marginal cost. It means that the cost of hiring worker is less than the benefits gained from them.